# Higher Lie characters and root enumeration in classical Weyl groups

### v1

@@ 7.1. Gelfand models
It follows from the Frobenius-Schur indicator theorem that for any real represented finite group $G$, $\varphi^{G}_{2}$ is a multiplicity-free sum of all irreducible $G$-characters, see [13, §4] and [21, Ex. 7.69]. Since $D_{n}$ is real represented, Theorem 1.6 completes the proof. ∎
@@ 7.2. Index 2 subgroups
Let $H$ be a subgroup of index $2$ of a classical Weyl group of type $A$, $B$ or $D$. For which integers $k$ is the root enumerator $\varphi^{H}_{k}$ a proper character?
@@
If $H$ is an index $2$ subgroup of a finite Coxeter group, then the square root enumerator $\varphi_{2}^{H}$ is a proper $H$-character.
@@
Every finite Coxeter group is totally orthogonal, i.e., all its irreducible representations are real; see, e.g., [16]. By [23, Proposition 3.6], if $H$ is an index $2$ subgroup of a totally orthogonal group then any real character of $H$ comes from a real representation. It follows, by the Frobenius-Schur indicator theorem (see [13, §4]), that $\varphi_{2}^{H}$ is a multiplicity-free sum of certain (not necessarily all) irreducible $H$-characters, and is therefore a proper character. ∎
@@
If $H$ is an index $2$ subgroup of a Weyl group of type $A$, $B$ or $D$, then $\varphi^{H}_{k}$ is proper for every odd $k$.
@@
If $n$ is odd then $A(B_{n})\cong D_{n}$, hence $\varphi^{A(B_{n})}_{k}$ is proper for any $k$.
@@
Hereby we list explicit generating functions for $\varphi_{k}^{H}$, where $H$ is ${\mathbb{Z}}_{2}\wr A(S_{n})$, $A(B_{n})$ or $A(D_{n})$. Recall, from Definition 6.1, that the linear character $\chi$ is defined on $B_{n}={\mathbb{Z}}_{2}\wr S_{n}$ as follows: $\chi$ is the identity map (i.e., the sign character) on ${\mathbb{Z}}_{2}=\{+1,-1\}$, and is trivial on the wreathing group $S_{n}$. Define now a linear character $\chi^{\prime}$ on $B_{n}$ as follows: $\chi^{\prime}$ is trivial on ${\mathbb{Z}}_{2}$, and is the sign character on the wreathing group $S_{n}$. For $n\geq 2$, the group $B_{n}$ has three nontrivial linear characters: $\chi$, $\chi^{\prime}$, and $\chi\chi^{\prime}$. They define the three index $2$ subgroups of $B_{n}$: $D_{n}=\ker(\chi)$, ${\mathbb{Z}}_{2}\wr A(S_{n})=\ker(\chi^{\prime})$, and $A(B_{n})=\ker(\chi\chi^{\prime})$.
@@
${\overline{\varphi}}_{k}^{B_{n}}(y):=\sum_{\begin{subarray}{c}x\in B_{n}\\ x^{k}=y\end{subarray}}\chi(x)\qquad(y\in B_{n}).$
@@
$\varphi_{k}^{\prime B_{n}}(y):=\sum_{\begin{subarray}{c}x\in B_{n}\\ x^{k}=y\end{subarray}}\chi^{\prime}(x)\qquad(y\in B_{n})$
@@
${\overline{\varphi}}_{k}^{\prime B_{n}}(y):=\sum_{\begin{subarray}{c}x\in B_{n}\\ x^{k}=y\end{subarray}}\chi\chi^{\prime}(x)\qquad(y\in B_{n}).$
@@
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\frac{\varphi_{k}^{\prime B_{n}}({\underline{\mu}})}{|Z_{\underline{\mu}}|}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}=\exp\left(\sum_{j,\theta}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}\frac{-(-t_{j,\theta})^{jk/h}}{2jk/h}\cdot\varphi_{h}^{{\mathbb{Z}}_{2}}(\theta)\right)$
@@
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\frac{{\overline{\varphi}}_{k}^{\prime B_{n}}({\underline{\mu}})}{|Z_{\underline{\mu}}|}{\underline{t}}^{{\underline{c}}({\underline{\mu}})}=\exp\left(\sum_{j,\theta}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}\frac{-(-t_{j,\theta})^{jk/h}}{2jk/h}\cdot{\overline{\varphi}}_{h}^{{\mathbb{Z}}_{2}}(\theta)\right),$
where $\varphi_{h}^{{\mathbb{Z}}_{2}}(\theta)=\sum_{\varepsilon\in{\mathbb{Z}}_{2},\,\varepsilon^{h}=\theta}1$ and ${\overline{\varphi}}_{h}^{{\mathbb{Z}}_{2}}(\theta)=\sum_{ve\in{\mathbb{Z}}_{2},\,\varepsilon^{h}=\theta}\chi(\varepsilon)$ $(h\in{\mathbb{Z}},\,\theta\in{\mathbb{Z}}_{2})$.
@@
The corresponding expression for the sign character of the symmetric group $S_{n}$ was obtained by Chernoff [5] and by Glebsky, Licón and Rivera [10]. In our notation, including the class function $\varphi_{k}^{\prime S_{n}}(y):=\sum_{x\in S_{n},\,x^{k}=y}\operatorname{sign}(x)$ (for $y\in S_{n}$), namely $\varphi_{k}^{\prime S_{n}}(\mu)$ (for a cycle type $\mu\vdash n$), their result is:
$\sum_{n\geq 0}\sum_{\mu\vdash n}\frac{\varphi_{k}^{\prime S_{n}}(\mu)}{|Z_{\mu}|}{\underline{t}}^{{\underline{c}}(\mu)}=\exp\left(\sum_{j\geq 1}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}\frac{-(-t_{j})^{jk/h}}{jk/h}\right).$
The following formulas are natural analogues of the corresponding formula for $\varphi_{k}^{D_{n}}$.
@@
$\varphi_{k}^{B_{n}}(y)+\varphi_{k}^{\prime B_{n}}(y)=\begin{cases}2\varphi_{k}^{{\mathbb{Z}}_{2}\wr A(S_{n})}(y),&\text{if }y\in{\mathbb{Z}}_{2}\wr A(S_{n});\\ 0,&\text{otherwise,}\end{cases}$
$\varphi_{k}^{B_{n}}(y)+{\overline{\varphi}}_{k}^{\prime B_{n}}(y)=\begin{cases}2\varphi_{k}^{A(B_{n})}(y),&\text{if }y\in A(B_{n});\\ 0,&\text{otherwise,}\end{cases}$
$\varphi_{k}^{B_{n}}(y)+{\overline{\varphi}}_{k}^{B_{n}}(y)+\varphi_{k}^{\prime B_{n}}(y)+{\overline{\varphi}}_{k}^{\prime B_{n}}(y)=\begin{cases}4\varphi_{k}^{A(D_{n})}(y),&\text{if }y\in A(D_{n});\\ 0,&\text{otherwise.}\end{cases}$
@@ 7.3. Quasisymmetric functions
In their seminal paper [9], Gessel and Reutenauer show that for every partition $\lambda\vdash n$, the Frobenius image of $\psi^{\lambda}$, the higher Lie character of type $A$, is equal to the sum of the fundamental quasisymmetric functions corresponding to the descent sets of all the permutations in the conjugacy class of cycle type $\lambda$.
@@
For every bipartition ${\underline{\lambda}}=(\lambda^{+},\lambda^{-})\vdash n$, the Frobenius characteristic image of the type $B$ higher Lie character $\psi^{\underline{\lambda}}$ is equal to the sum of Poirier’s fundamental quasisymmetric functions corresponding to the signed descent sets of all the elements in the conjugacy class of signed cycle type ${\underline{\lambda}}$.

### v2

@@ 7.1. Gelfand models
It follows from the Frobenius-Schur indicator theorem that for any real represented finite group $G$, $\rho^{G}_{2}$ is a multiplicity-free sum of all irreducible $G$-characters, see [13, §4] and [21, Ex. 7.69]. Since $D_{n}$ is real represented, Theorem 1.6 completes the proof. ∎
@@ 7.2. Index 2 subgroups
Let $H$ be a subgroup of index $2$ of a classical Weyl group of type $A$, $B$ or $D$. For which integers $k$ is the root enumerator $\rho^{H}_{k}$ a proper character?
@@
If $H$ is an index $2$ subgroup of a finite Coxeter group, then the square root enumerator $\rho_{2}^{H}$ is a proper $H$-character.
@@
Every finite Coxeter group is totally orthogonal, i.e., all its irreducible representations are real; see, e.g., [16]. By [23, Proposition 3.6], if $H$ is an index $2$ subgroup of a totally orthogonal group then any real character of $H$ comes from a real representation. It follows, by the Frobenius-Schur indicator theorem (see [13, §4]), that $\rho_{2}^{H}$ is a multiplicity-free sum of certain (not necessarily all) irreducible $H$-characters, and is therefore a proper character. ∎
@@
If $H$ is an index $2$ subgroup of a Weyl group of type $A$, $B$ or $D$, then $\rho^{H}_{k}$ is proper for every odd $k$.
@@
If $n$ is odd then $A(B_{n})\cong D_{n}$, hence $\rho^{A(B_{n})}_{k}$ is proper for any $k$.
@@
Hereby we list explicit generating functions for $\rho_{k}^{H}$, where $H$ is ${\mathbb{Z}}_{2}\wr A(S_{n})$, $A(B_{n})$ or $A(D_{n})$. Recall, from Definition 6.1, that the linear character $\chi$ is defined on $B_{n}={\mathbb{Z}}_{2}\wr S_{n}$ as follows: $\chi$ is the identity map (i.e., the sign character) on ${\mathbb{Z}}_{2}=\{+1,-1\}$, and is trivial on the wreathing group $S_{n}$. Define now a linear character $\chi^{\prime}$ on $B_{n}$ as follows: $\chi^{\prime}$ is trivial on ${\mathbb{Z}}_{2}$, and is the sign character on the wreathing group $S_{n}$. For $n\geq 2$, the group $B_{n}$ has three nontrivial linear characters: $\chi$, $\chi^{\prime}$, and $\chi\chi^{\prime}$. They define the three index $2$ subgroups of $B_{n}$: $D_{n}=\ker(\chi)$, ${\mathbb{Z}}_{2}\wr A(S_{n})=\ker(\chi^{\prime})$, and $A(B_{n})=\ker(\chi\chi^{\prime})$.
@@
${\overline{\rho}}_{k}^{B_{n}}(y):=\sum_{\begin{subarray}{c}x\in B_{n}\\ x^{k}=y\end{subarray}}\chi(x)\qquad(y\in B_{n}).$
@@
$\rho_{k}^{\prime B_{n}}(y):=\sum_{\begin{subarray}{c}x\in B_{n}\\ x^{k}=y\end{subarray}}\chi^{\prime}(x)\qquad(y\in B_{n})$
@@
${\overline{\rho}}_{k}^{\prime B_{n}}(y):=\sum_{\begin{subarray}{c}x\in B_{n}\\ x^{k}=y\end{subarray}}\chi\chi^{\prime}(x)\qquad(y\in B_{n}).$
@@
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}\rho_{k}^{\prime B_{n}}({\underline{\mu}})\,\frac{{\underline{t}}^{{\underline{c}}({\underline{\mu}})}}{|Z_{\underline{\mu}}|}=\exp\left(\sum_{j,\theta}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}\rho_{h}^{{\mathbb{Z}}_{2}}(\theta)\,\frac{-(-t_{j,\theta})^{jk/h}}{2jk/h}\right)$
@@
$\sum_{n\geq 0}\sum_{{\underline{\mu}}\vdash n}{\overline{\rho}}_{k}^{\prime B_{n}}({\underline{\mu}})\,\frac{{\underline{t}}^{{\underline{c}}({\underline{\mu}})}}{|Z_{\underline{\mu}}|}=\exp\left(\sum_{j,\theta}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}{\overline{\rho}}_{h}^{{\mathbb{Z}}_{2}}(\theta)\,\frac{-(-t_{j,\theta})^{jk/h}}{2jk/h}\right),$
where $\rho_{h}^{{\mathbb{Z}}_{2}}(\theta)=\sum_{\varepsilon\in{\mathbb{Z}}_{2},\,\varepsilon^{h}=\theta}1$ and ${\overline{\rho}}_{h}^{{\mathbb{Z}}_{2}}(\theta)=\sum_{ve\in{\mathbb{Z}}_{2},\,\varepsilon^{h}=\theta}\chi(\varepsilon)$ $(h\in{\mathbb{Z}},\,\theta\in{\mathbb{Z}}_{2})$.
@@
The corresponding expression for the sign character of the symmetric group $S_{n}$ was obtained by Chernoff [5] and by Glebsky, Licón and Rivera [10]. In our notation, including the class function $\rho_{k}^{\prime S_{n}}(y):=\sum_{x\in S_{n},\,x^{k}=y}\operatorname{sign}(x)$ (for $y\in S_{n}$), namely $\rho_{k}^{\prime S_{n}}(\mu)$ (for a cycle type $\mu\vdash n$), their result is:
$\sum_{n\geq 0}\sum_{\mu\vdash n}\rho_{k}^{\prime S_{n}}(\mu)\,\frac{{\underline{t}}^{{\underline{c}}(\mu)}}{|Z_{\mu}|}=\exp\left(\sum_{j\geq 1}\sum_{\begin{subarray}{c}h|k\\ \gcd(h,j)=1\end{subarray}}\frac{-(-t_{j})^{jk/h}}{jk/h}\right).$
The following formulas are natural analogues of the corresponding formula for $\rho_{k}^{D_{n}}$.
@@
$\rho_{k}^{B_{n}}(y)+\rho_{k}^{\prime B_{n}}(y)=\begin{cases}2\rho_{k}^{{\mathbb{Z}}_{2}\wr A(S_{n})}(y),&\text{if }y\in{\mathbb{Z}}_{2}\wr A(S_{n});\\ 0,&\text{otherwise,}\end{cases}$
$\rho_{k}^{B_{n}}(y)+{\overline{\rho}}_{k}^{\prime B_{n}}(y)=\begin{cases}2\rho_{k}^{A(B_{n})}(y),&\text{if }y\in A(B_{n});\\ 0,&\text{otherwise,}\end{cases}$
$\rho_{k}^{B_{n}}(y)+{\overline{\rho}}_{k}^{B_{n}}(y)+\rho_{k}^{\prime B_{n}}(y)+{\overline{\rho}}_{k}^{\prime B_{n}}(y)=\begin{cases}4\rho_{k}^{A(D_{n})}(y),&\text{if }y\in A(D_{n});\\ 0,&\text{otherwise.}\end{cases}$
@@ 7.3. Quasisymmetric functions
In their seminal paper [9], Gessel and Reutenauer show that for every partition $\lambda\vdash n$, the Frobenius image of $\psi_{S_{n}}^{\lambda}$, the higher Lie character of type $A$, is equal to the sum of the fundamental quasisymmetric functions corresponding to the descent sets of all the permutations in the conjugacy class of cycle type $\lambda$.
@@
For every bipartition ${\underline{\lambda}}=(\lambda^{+},\lambda^{-})\vdash n$, the Frobenius characteristic image of the type $B$ higher Lie character $\psi_{B_{n}}^{\underline{\lambda}}$ is equal to the sum of Poirier’s fundamental quasisymmetric functions corresponding to the signed descent sets of all the elements in the conjugacy class of signed cycle type ${\underline{\lambda}}$.